# Quasinormal modes of Schwarzschild-de Sitter black holes in semi-open systems

### v2

@@ I Introduction
Black holes are non-Hermitian (NH) systems. Such NH systems exist in many fields of physics Ashida et al. (2021). A hallmark of NH systems is the existence of exceptional point (EP) Heiss (1999, 2012); Ryu et al. (2025); Zhong et al. (2018); Dietz et al. (2011), defined as a spectral singularity where both eigenvalues and their corresponding eigenvectors coalesce. For more mathematical aspects of EPs, we can refer to monograph Kato (1976). For the gravity aspects, EPs also appear in the QNM framework of black hole perturbation systems. For massive scalar perturbation of a Kerr black hole, two QNM spectra can become degenerate Cavalcante et al. (2024a, b, 2025), where the EP is $(M\mu)_{c}\simeq 0.3704981$ and $(a/M)_{c}\simeq 0.9994660$. Such EP is also obtained in Chen et al. (2025a). The EPs can also be found when a Gaussian bump is added to the Regge-Wheeler effective potential Yang et al. (2025a), where the EP is $\epsilon\simeq 0.005$, and $d\simeq 15.698$. Near such EPs, the conventional decomposition of the ringdown waveform into a superposition of QNMs may become invalid Yang et al. (2025a); Panosso Macedo et al. (2025). A phenomenon related to the concept of exceptional points is called mode repulsion or avoided crossing, which have been widely studied in Dias et al. (2022); Motohashi (2025); Oshita et al. (2025); Lo et al. (2025); Takahashi et al. (2025). When the dimension of parameters increases, EPs will form exceptional lines (ELs) which widely arises in non-Hermitian systems Shen et al. (2018); Yang and Hu (2019); Wu et al. (2024). For the gravity aspect, in Cao et al. (2025a), EL is found, and the topological structures (such Berry phase and vorticity) of QNMs associated with EL are also considered. Although we know that the spectrum is unstable, the spectra near exceptional points will be even more unstable, which is studied through the pseudospectra at exceptional points in Cao et al. (2025a).
@@ IV Greybody factors
Figure 3: Greybody factor $\Gamma(\omega)$ for axial gravitational $(s=2)$ perturbations of the Schwarzschild-de Sitter black hole with a frequency-independent reflectivity $\mathcal{K}(\omega)$. In both panels the reflective wall is placed at $x_{0}=-30$ ($\Delta=4.5730021\times 10^{-3}$ and $\Delta=2.1169523\times 10^{-12}$), and the curves with different colors correspond to $\mathcal{K}(\omega)=0,\,0.25,\,0.50,\,0.75,$ and $1$. Panel (a) shows the near-extremal case with $r_{e}=1$ and $r_{c}=1.1$, and panel (b) shows the nearly Schwarzschild case with $r_{e}=1$ and $r_{c}=100$.
@@ VI Conclusions and discussion
In the theoretical setup, we begin by reviewing the geometry of the SdS spacetime and cast the perturbation equation into the form of Heun’s equation (cf. Appendix A). Local analytic solutions near the event and cosmological horizons are obtained (see Eq. (6) and Eq. (9)). Connection coefficients between solutions in different regions are established using Wronskians. The boundary conditions for present semi-open system are precisely defined as: at a fixed position near the event horizon, the wave function is a superposition of ingoing and outgoing waves, and at the cosmological horizon, only outgoing waves are required. Hence, the resonance condition determining the QNM spectra is derived, which is written as $S_{\text{in}}=0$.
Regarding the QNM spectra (see Sec. III), we focuses on axial gravitational perturbations $(s=2)$ and examine the impact of a real, constant reflectivity, i.e., $\mathcal{K}\in[0,1]$. Two paradigmatic cases are compared, the near-extremal SdS black hole and the nearly-Schwarzschild case. Numerical computations reveal that as the reflectivity increases from 0 (open system) to $1$ (perfectly reflective wall), the QNMs undergo systematic migration in the complex plane. Three distinct types of migrations are identified: a subset of modes approaches the real axis, forming long-lived QBS; another subset moves toward but does not reach the real axis retaining a non-zero decay rate; the third type, primarily higher overtones, veers toward and eventually lands on the imaginary axis, becoming purely decaying modes. It is also found that the position of the reflective wall affects spectrum stability. To be more specific, walls placed closer to the event horizon render the spectrum more sensitive to small changes in $\mathcal{K}$, indicating enhanced instability.
@@ Appendix A The Heun’s equation
These four points happen to be the four singular points of the Heun equation. We next perform the scaling transformation of $\Psi(z)$:
@@
Next, we give parameters in the Heun equation for all cases, where the condition $r_{n}=-(r_{e}+r_{c})$ has been used. Note that $\alpha$ and $\beta$ can be exchanged. For the three types of perturbations, they share the same $\gamma$, $\delta$, and $\epsilon$, namely,
@@
$\mathrm{e}^{\mathrm{i}\omega x}$ $=$ $\Big(1-\frac{r_{e}}{r_{c}}\Big)^{\rho_{e,1}}\Big(\frac{r_{c}}{r_{e}}-1\Big)^{\rho_{c,1}}\Big(\frac{r_{n}}{r_{e}}-\frac{r_{n}}{r_{c}}\Big)^{\rho_{n,1}}z^{\rho_{e,1}}(1-z)^{\rho_{c,1}}(z-z_{n})^{\rho_{n,1}}\,$,
@@
Combining Eq. (48) with Eqs. (B), one solves $B_{\text{out}}(\omega)$ and $B_{\text{in}}(\omega)$ which are given by
$B_{\text{in}}(\omega)=1/F(\omega)\cdot C_{12}(\omega)\,,\quad B_{\text{out}}(\omega)=C_{11}(\omega)\,$, (49)
@@
$\Psi_{\text{in}}=Y_{\text{in}}(\omega)\cdot\Psi_{02}\,,\quad\Psi_{\text{up}}=Y_{\text{out}}(\omega)\cdot\Psi_{11}\,$, (54)
in which two factor functions $Y_{\text{in}}(\omega)$ and $Y_{\text{out}}(\omega)$ are
@@
$Y_{\text{out}}(\omega)=\Big(1-\frac{r_{e}}{r_{c}}\Big)^{\rho_{e,1}}\Big(\frac{r_{c}}{r_{e}}-1\Big)^{\rho_{c,1}}\Big(\frac{r_{n}}{r_{e}}-\frac{r_{n}}{r_{c}}\Big)^{\rho_{n,1}}(-1)^{\rho_{c,2}}\,.$ (56)

### v3

@@ I Introduction
Black holes are NH systems. Such NH systems exist in many fields of physics Ashida et al. (2021). A hallmark of NH systems is the existence of exceptional point (EP) Heiss (1999, 2012); Ryu et al. (2025); Zhong et al. (2018); Dietz et al. (2011), defined as a spectral singularity where both eigenvalues and their corresponding eigenvectors coalesce. For more mathematical aspects of EPs, we can refer to monograph Kato (1976). For the gravity aspects, EPs also appear in the QNM framework of black hole perturbation systems. For massive scalar perturbation of a Kerr black hole, two QNM spectra can become degenerate Cavalcante et al. (2024a, b, 2025), where the EP is $(M\mu)_{c}\simeq 0.3704981$ and $(a/M)_{c}\simeq 0.9994660$. Such EP is also obtained in Chen et al. (2025a). The EPs can also be found when a Gaussian bump is added to the Regge-Wheeler effective potential Yang et al. (2025a), where the EP is $\epsilon\simeq 0.005$, and $d\simeq 15.698$. Near such EPs, the conventional decomposition of the ringdown waveform into a superposition of QNMs may become invalid Yang et al. (2025a); Panosso Macedo et al. (2025). A phenomenon related to the concept of exceptional points is called mode repulsion or avoided crossing, which have been widely studied in Dias et al. (2022); Motohashi (2025); Oshita et al. (2025); Lo et al. (2025); Takahashi et al. (2025). When the dimension of parameters increases, EPs will form exceptional lines (ELs) which widely arises in non-Hermitian systems Shen et al. (2018); Yang and Hu (2019); Wu et al. (2024). For the gravity aspect, in Cao et al. (2025a), EL is found, and the topological structures (such Berry phase and vorticity) of QNMs associated with EL are also considered. Although we know that the spectrum is unstable, the spectra near exceptional points will be even more unstable, which is studied through the pseudospectra at exceptional points in Cao et al. (2025a).
@@ IV Greybody factors
Figure 3: Greybody factor $\Gamma(\omega)$ for axial gravitational $(s=2)$ perturbations of the Schwarzschild-de Sitter black hole with a frequency-independent reflectivity $\mathcal{K}$. In both panels the reflective wall is placed at $x_{0}=-30$ ($\Delta=4.5730021\times 10^{-3}$ and $\Delta=2.1169523\times 10^{-12}$), and the curves with different colors correspond to $\mathcal{K}=0,\,0.25,\,0.50,\,0.75,$ and $1$. Panel (a) shows the near-extremal case with $r_{e}=1$ and $r_{c}=1.1$, and panel (b) shows the near Schwarzschild case with $r_{e}=1$ and $r_{c}=100$.
@@ VI Conclusions and discussion
In the theoretical setup, we begin by reviewing the geometry of the SdS spacetime and cast the perturbation equation into the form of Heun’s equation (cf. Appendix A). Local analytic solutions near the event and cosmological horizons are obtained (see Eq. (6) and Eq. (9)). Connection coefficients between solutions in different regions are established using Wronskians. The boundary conditions for present semi-open system are precisely defined as: at a fixed position near the event horizon, the wave function is a superposition of ingoing and outgoing waves, and at the cosmological horizon, only outgoing waves are required. Hence, the resonance condition determining the QNM spectra is derived, which is written as $S_{\text{in}}(\omega)=0$.
Regarding the QNM spectra (see Sec. III), we focuses on axial gravitational perturbations $(s=2)$ and examine the impact of a real, constant reflectivity, i.e., $\mathcal{K}\in[0,1]$. Two paradigmatic cases are compared, the near-extremal SdS black hole and the near Schwarzschild case. Numerical computations reveal that as the reflectivity increases from 0 (open system) to $1$ (perfectly reflective wall), the QNMs undergo systematic migration in the complex plane. Three distinct types of migrations are identified: a subset of modes approaches the real axis, forming long-lived QBS; another subset moves toward but does not reach the real axis retaining a non-zero decay rate; the third type, primarily higher overtones, veers toward and eventually lands on the imaginary axis, becoming purely decaying modes. It is also found that the position of the reflective wall affects spectrum stability. To be more specific, walls placed closer to the event horizon render the spectrum more sensitive to small changes in $\mathcal{K}$, indicating enhanced instability.
@@ Appendix A The Heun’s equation
These four points happen to be the four regular singular points of the Heun’s equation. We next perform the scaling transformation of $\Psi(z)$:
@@
Next, we give parameters in the Heun’s equation for all cases, where the condition $r_{n}=-(r_{e}+r_{c})$ has been used. Note that $\alpha$ and $\beta$ can be exchanged. For the three types of perturbations, they share the same $\gamma$, $\delta$, and $\epsilon$, namely,
@@
$\mathrm{e}^{\mathrm{i}\omega x}$ $=$ $\Big(1-\frac{r_{e}}{r_{c}}\Big)^{\rho_{e,1}}\Big(\frac{r_{c}}{r_{e}}-1\Big)^{\rho_{c,1}}\Big(\frac{r_{n}}{r_{e}}-\frac{r_{n}}{r_{c}}\Big)^{\rho_{n,1}}z^{\rho_{e,1}}(1-z)^{\rho_{c,1}}(z-z_{n})^{\rho_{n,1}}\,$, (47)
@@
Combining Eq. (49) with Eqs. (47), one solves $B_{\text{out}}(\omega)$ and $B_{\text{in}}(\omega)$ which are given by
$B_{\text{in}}(\omega)=C_{12}(\omega)/F(\omega)\,,\quad B_{\text{out}}(\omega)=C_{11}(\omega)\,$, (50)
@@
$\Psi_{\text{in}}=Y_{\text{in}}(\omega)\cdot\Psi_{02}\,,\quad\Psi_{\text{up}}=Y_{\text{up}}(\omega)\cdot\Psi_{11}\,$, (55)
in which two factor functions $Y_{\text{in}}(\omega)$ and $Y_{\text{up}}(\omega)$ are
@@
$Y_{\text{up}}(\omega)=\Big(1-\frac{r_{e}}{r_{c}}\Big)^{\rho_{e,1}}\Big(\frac{r_{c}}{r_{e}}-1\Big)^{\rho_{c,1}}\Big(\frac{r_{n}}{r_{e}}-\frac{r_{n}}{r_{c}}\Big)^{\rho_{n,1}}(-1)^{\rho_{c,2}}\,.$ (57)